A cantilever of length 2 m carries two point loads of 20 KN at the free end and 30 KN at a distance of 1 m from the free end. Find the slope and deflection at free end. Take E= 2×10^5 N/mm^2 and I= 1.5×10^8 mm^4

1
Expert's answer
2022-02-02T12:12:36-0500

Given quantities:

"l = 2m"

"E=2*10^5 \frac{N}{mm^2}"

"I=1.6*10^8 mm^4"

"\theta-?"

"\theta = \frac{wl^2}{2EI}=\frac{20*2^2}{2*12*10^2}=0.0033rad"